OpenStudy (anonymous):

What is the slope of the line that contains the points (-8, 7) and (-4, -5)? A. -3 B.1/3 C.3 D.-1/3

OpenStudy (anonymous):

the answer is [y2-y1\div x2-x1\]...ie.....y2-y1=-5-7=-12 and x2-x1=-4-(-8)=+4 hence the answer is -12/4=-3 THE ANS IS ----(A)=-3

OpenStudy (anonymous):

using the formula \[{y_{2} - y_{1}}\over {x_{2}-x_{1}}\]we plug in the values \[{-5-7}\over-4-(-8)\]\[{-5-7\over-4+8}\]combine like terms\[-12\over4\]simplify\[=-3\]so your slope is -3 :)